# ASGO: Adaptive Structured Gradient Optimization

### v1

@@ 6 More Discussions on the Algorithm
Figure 2: This figure is from Zhang et al. (2024b), describing the Hessian of different parameter blocks in a small Transformer at 1% training step. The near-block-diagonal structure maintains throughout training. But different parameter blocks have different numbers of small dense matrices, where Query and Key correspond to the number of heads.
Special Design for Transformers.
Furthermore, we introduce a specialized adaptation of ASGO for query and key matrices in attention layers. Recent work by Zhang et al. (2024b) has demonstrated that the Hessian structure of query and key layers differs significantly from conventional MLP layers. As illustrated in Figure 2, the number of dense blocks corresponds to the number of attention heads rather than output neurons. This observation aligns with the forward computation of multi-head attention, where attention scores are computed independently across different subspaces, suggesting that query and key parameters could be optimized in a head-wise manner. To leverage this insight, we reshape query and key parameters from matrices in $R^{n\times hd}$ to three-dimensional tensors $R^{h\times n\times d}$, and apply our optimization algorithm independently to each head’s subspace. This restructuring reduces both memory consumption and computational complexity by decreasing the matrix size from $O(h^{2}d^{2})$ into $O(hd^{2})$. For instance, in NanoGPT (Karpathy, 2022), this adjustment reduces the preconditioning size of a single query/key parameter from approximately $10^{6}$ elements into $10^{5}$. The empirical benefits of this modification are evaluated in Section 7.
A Diagonal Variant of ASGO.
Drawing inspiration from Duchi et al. (2011), we also implement a variant of ASGO using diagonal matrices, which we denote as DASGO in Algorithm 3. DASGO can be viewed as a light version of ASGO, which eliminates the need to compute the square root and inverse of matrices and reduces the memory requirements to a comparable level with Adam-mini (Zhang et al., 2024b). However, since DASGO only employs a diagonal preconditioner, it fails to recover the superior theoretical properties of ASGO under block-wise diagonal Hessian settings. We further empirically examine this tradeoff in the next section.
@@
1:  Input: $W_{0}\in\mathbb{R}^{m\times n}$, $\epsilon\in\mathbb{R}$, schedule $\{\eta_{t}\}$, momentums $\beta_{1},\beta_{2}\in[0,1)$, and batch size $M\in\mathbb{N}$, update interval $\tau\in\mathbb{N}$ ($\epsilon$ should be small, similar to the $\epsilon$ for Adam or AdaGrad)
2:  Initialize $M_{-1}=0\in\mathbb{R}^{m\times n}$, $V_{-1}=0\in\mathbb{R}^{n\times n}$
@@
5:     $G_{t}=\frac{1}{M}\sum_{\xi\in\mathcal{B}_{t}}\nabla_{W}f(W_{t};\xi)$triangleright Compute stochastic gradient
6:     $M_{t}=\beta_{1}M_{t-1}+(1-\beta_{1})G_{t}$triangleright Accumulate momentum
7:     $V_{t}=\beta_{2}V_{t-1}+(1-\beta_{2})G_{t}^{\top}G_{t}$triangleright Use exponential moving average to update $V_{t}$
8:     if $\mathrm{mod}(t,\tau)=0$triangleright Update $\Lambda_{t}^{\mathrm{inv}}$ every $\tau$ iterations
9:      $\Lambda_{t}^{\mathrm{inv}}=\left(V_{t}+\epsilon I_{n}\right)^{-\frac{1}{2}}$
10:     else
11:      $\Lambda_{t}^{\mathrm{inv}}=\Lambda_{t-1}^{\mathrm{inv}}$
12:     end if
13:     $W_{t+1}=W_{t}-\eta_{t}M_{t}\Lambda_{t}^{\mathrm{inv}}$triangleright Precondition from the right side
14:  end for
@@
1:  Input: $W_{0}\in\mathbb{R}^{m\times n}$, $\epsilon\in\mathbb{R}$, schedule $\{\eta_{t}\}$, momentums $\beta_{1},\beta_{2}\in[0,1)$, and batch size $M\in\mathbb{N}$
@@
5:     $G_{t}=\frac{1}{M}\sum_{\xi\in\mathcal{B}_{t}}\nabla_{W}f(W_{t};\xi)$triangleright Compute stochastic gradient
6:     $M_{t}=\beta_{1}M_{t-1}+(1-\beta_{1})G_{t}$triangleright Accumulate momentum
7:     $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})\operatorname{diag}(G_{t}^{\top}G_{t}$) $\triangleright$ Use exponential moving average to update vector $V_{t}$
8:     $W_{t+1}=W_{t}-\eta_{t}M_{t}\operatorname{diag}(v_{t}+\epsilon)^{-\frac{1}{2}}$triangleright Precondition by a diagonal Matrix from right side
@@
7 Empirical Results
We empirically evaluate the effectiveness of ASGO (Algorithm 2) and DASGO (Algorithm 3) on pretraining tasks for deep learning models. We compare the methods against established optimizers, including AdamW (Kingma and Ba, 2014; Loshchilov and Hutter, 2017), Shampoo (Gupta et al., 2018), and Muon (Jordan et al., 2024). Several important implementation details should be noted for fair comparison. First, since Muon is designed to operate exclusively on matrix and tensor parameters, we follow Jordan et al. (2024) to apply AdamW update rules to all 1D parameters within the Muon optimizer to ensure that it can handle the complete model. Additionally, the specialized attention layer modification for query and key matrices described in Section 6 can be applied to Muon as well, providing similar benefits in reducing memory consumption and computational complexity.
Recent research (Shi et al., 2023; Morwani et al., 2024; Lin et al., 2024) has suggested treating the inverse order in Shampoo as a tunable hyperparameter rather than using the standard value of $1/4$. Notably, Morwani et al. (2024) demonstrated that Shampoo2 (Shampoo with inverse-order $1/2$) provides a superior approximation of the full-matrix AdaGrad optimizer. Motivated by this finding, we treat the inverse order as a hyperparameter in our pretraining experiments. Additionally, Shampoo’s performance is highly dependent on the initialization of its preconditioner matrices, as a well-initialized preconditioner is crucial for accurately approximating the Empirical Fisher Information Matrix (Morwani et al., 2024). To address this initialization sensitivity, we implement a preconditioner warmup phase specifically for the Shampoo optimizer. Following the methodology in Ren and Goldfarb (2021), we dedicate the first epoch exclusively to accumulating statistics for the preconditioner matrices without updating model parameters. This approach yields a more robust estimation of the preconditioner, which significantly improves the stability of the Shampoo algorithm during subsequent training iterations.
@@ 7 Empirical Results
7.1 Pretraining NanoGPT
We conducted experiments using the NanoGPT architecture on the Shakespeare character dataset, following the configuration in (Karpathy, 2022). To ensure a fair comparison, we maintained consistent hyperparameter (HP) settings across all optimization methods, varying only those parameters that significantly impacted optimizer performance, i.e., learning rate, 1st and 2nd order moment coefficients ($\beta_{1}$ and $\beta_{2}$), preconditioner update frequency ($\tau$), and warm-up ratio  [^3] A full description of how these “best” HPs were selected and their values is presented in Appendix A.. Figure 3 presents a comparison of the pretraining performance of ASGO and DASGO with that of Shampoo, Muon and AdamW on the NanoGPT model. Examining both train and test loss curves, ASGO consistently outperforms Shampoo despite requiring only half the memory consumption and computational complexity, which highlights ASGO’s practical advantages for training language models. Furthermore, ASGO and Muon achieve the lowest final training and test losses, outperforming all other methods. This consistent performance between ASGO and Muon aligns well with our discussions in Section 6.
The lightweight DASGO optimizer achieves competitive results against AdamW in both training and test loss metrics. DASGO demonstrates particularly strong performance during the initial training phase (first 200 training step), where it converges faster than other methods. This early convergence advantage suggests that DASGO could serve as an effective warmup optimizer in multi-stage training pipelines. On the other hand, ASGO demonstrates more stable convergence with less oscillation in later training stages compared to DASGO. This performance gap between DASGO and ASGO highlights the value of preserving non-diagonal elements in the preconditioner matrices, particularly for capturing parameter interactions in attention-based architectures. Nevertheless, DASGO’s competitive performance while requiring substantially fewer computational resources makes it an attractive option for resource-constrained scenarios.
@@ 7.1 Pretraining NanoGPT
Ablation Study: Impact of Special Query/Key Design
In Section 6, we highlighted the computational benefits of specialized processing for Query and Key matrices in Transformer architectures. To validate these theoretical advantages empirically, we conducted an ablation study comparing the performance and training stability of ASGO and Muon, both with and without the specialized Query-Key processing. We maintained the optimal hyperparameter configuration established in Section 7.1 for all parameters except the learning rate. For each algorithm variant (with and without specialized processing), we randomly sampled learning rates from a log-uniform distribution ranging from $10^{-5}$ to $10^{-1}$. Each configuration was trained for 5 epochs on the NanoGPT model, after which we recorded the validation loss. This approach allowed us to evaluate both performance and robustness to learning rate selection. Figure 4 presents the distribution of validation losses after 5 epochs for both algorithms. The violin plots [^4] Violin plot outlines depict empirical probability density; i.e., the width of the shaded area represents the proportion of the data located there. Box plots within a violin plot display the median and inter-quartile range. reveal several key insights: (1) For ASGO (Figure 4a), the benefits of specialized Query-Key processing are substantial. The specialized implementation demonstrates a narrower, lower distribution of validation losses (centered around 2.5), indicating greater stability across different learning rates. In contrast, the vanilla implementation shows a long-tailed distribution extending beyond loss values of 6.0, with several outliers representing training instability at certain learning rates. (2) For Muon (Figure 4b), both implementations demonstrate nearly identical performance distributions, with their median validation losses and distribution shapes showing minimal differences. However, the specialized implementation achieves this comparable performance while reducing memory requirements and computational complexity. This represents a clear efficiency advantage—the specialized Query-Key processing effectively provides computational savings with no performance penalty.
(a) ASGO
(b) Muon
8 Conclusions
In this paper, we propose a novel algorithm ASGO, aiming to utilize the structured properties of optimization problems. We prove that ASGO achieves significantly better convergence rates compared to full-matrix AdaGrad and Shampoo. Based on the theory, we demonstrate that ASGO can benefit from low-rank gradients and block-wise diagonal Hessians, which are widely observed structured properties of deep neural networks. We further propose some practical modifications on ASGO, and empirical results on language model tasks verify the effectiveness of our approach. For future work, we plan to extend our research in two key directions:
  • Larger Foundation Models: We aim to evaluate the algorithm’s performance on large foundation models to demonstrate its effectiveness in pretraining LLMs.
  • We intend to generalize our algorithm to handle tensor cases, which would significantly expand its applicability across a broader range of neural network architectures.
Appendix A Details of Empirical Experiments
For our experiments in Section 7.1, we employ the NanoGPT architecture as given in Karpathy [2022]. The model consists of 6 Transformer layers, 6 attention heads, and an embedding dimension of 384. We utilize the OneCycle learning rate schedule consistently across all optimization algorithms to ensure fair comparison. Training is conducted on the Shakespeare character-level dataset with a sequence length of 256 tokens.
To ensure optimal performance for each optimizer, we conduct a comprehensive hyperparameter search using random search  [Bergstra and Bengio, 2012, Choi et al., 2019] within the following predefined ranges: learning rate(lr): $[10^{-5},10^{-1}]$; $\beta_{1}$: $[0.7,0.99]$; $\beta_{2}$: $[0.7,0.99]$; and select the warmup factor from the set $[0.1,0.15,0.2,0.25,0.3]$. For the optimizers Shampoo and ASGO, we select the update frequency $\tau$ from the list $[5,10,15]$. We also treat the inverse order in Shampoo as a hyperparameter and select it as either $-\frac{1}{2}$ or $-\frac{1}{4}$. HPs were selected based on validation loss after 20 epochs of training. For each optimizer, we performed a random search for 16 hours to find the optimal HP combination. These best HP choices for training NanoGPT are given in Table 1.

### v2

@@ Appendix A Additional Related Work
Figure 3: This figure is from Zhang et al. [2024b]. It depicts the Hessian of different parameter blocks in a small Transformer at the 1% training step. The near-block-diagonal structure maintains throughout training. But different parameter blocks have different numbers of small dense matrices, where Query and Key correspond to the number of heads.
Block-wise Diagonal Hessian.
It has been observed that the Hessian of a neural network tends to be block-wise diagonal with each block corresponding to a neuron both in experiments and theory for small MLPs [Collobert, 2004]. Zhang et al. [2024a, b] recently numerically verified this property in small transformers, and as illustrated in Figure 4, further empirically showed that transformers may exhibit heterogeneity between blocks, while CNNs may not.
(a) VGG16
(b) ViT-base
(c) GPT2
Figure 4: This figure is from Zhang et al. [2024a, b], calculating the Jensen-Shannon (JS) distance between two eigenvalue densities of all possible block pairs at initialization. It shows that JS distance of blockwise spectra in CNNs is significantly smaller than that in Transformers.
Concurrent Work.
When we were finishing writing this paper, we noticed that the paper [Xie et al., 2025], which had just appeared on arXiv, proposes and studies an algorithm, referred to as One-Sided Shampoo, that is identical to ASGO. Although the theoretical techniques and convergence results proved in  [Xie et al., 2025] and here are very similar in general, there are still some notable differences between the two works. First, the motivations are different. In Xie et al. [2025], the authors develop their method from the unified preconditioning method framework AdaptReg [Gupta et al., 2017] and mainly highlight its superior convergence results compared to Shampoo and full-matrix AdaGrad. In our paper, we focus more on theoretically discussing how ASGO can utilize the structured properties of optimization problems including low-rank gradients and block-wise diagonal Hessians, to highlight the potential of ASGO as a practical algorithm for training deep learning models. Second, our empirical results provide evidence that ASGO can perform well in practical tasks, whereas Xie et al. [2025] focuses on convex settings and examines One-Sided Shampoo only on linear regression. Moreover, our implementation includes specialized designs for Transformer architectures (particularly query/key) to improve performance on deep learning training tasks. As a direct byproduct of our main algorithm, we also developed a lightweight diagonal version named DASGO to trade off memory consumption and performance. Third, Xie et al. [2025] focuses primarily on developing a general proof framework applicable to multiple optimizers including One-Sided Shampoo, whereas our work specifically examines the theoretical and practical benefits of ASGO. To conclude, both works contribute to a better understanding of ASGO/One-Sided Shampoo.
Appendix B More Discussions on ASGO
A practical implementation of ASGO.
We present a practical implementation of ASGO in Algorithm 2. Our implementation incorporates several common modifications to enhance computational efficiency and stability.
@@ Appendix B More Discussions on ASGO
1:  Input: $W_{0}\in\mathbb{R}^{m\times n}$, lr schedule $\{\eta_{t}\}$, momentums $\beta_{1},\beta_{2}\in[0,1)$, batch size $M\in\mathbb{N}$, update interval $\tau\in\mathbb{N}$, $\epsilon\in\mathbb{R}$ ($\epsilon$ should be small, similar to the $\epsilon$ for Adam or AdaGrad.)
2:  Initialize $M_{-1}=0\in\mathbb{R}^{m\times n}$, $V_{-1}=0\in\mathbb{R}^{n\times n}\text{ or }\mathbb{R}^{m\times m}$
@@
5:     $G_{t}=\frac{1}{M}\sum_{\xi\in\mathcal{B}_{t}}\nabla_{W}f(W_{t};\xi)$
6:     $M_{t}=\beta_{1}M_{t-1}+(1-\beta_{1})G_{t}$
7:     if $m<n$ then
8:        $V_{t}=\beta_{2}V_{t-1}+(1-\beta_{2})G_{t}G_{t}^{\top}$
9:     else
10:        $V_{t}=\beta_{2}V_{t-1}+(1-\beta_{2})G_{t}^{\top}G_{t}$
11:     end if
12:     if $\mathrm{mod}(t,\tau)=0$
13:      $\Lambda_{t}^{\mathrm{inv}}=\left(V_{t}+\epsilon I_{n}\right)^{-\frac{1}{2}}$
14:     else
15:      $\Lambda_{t}^{\mathrm{inv}}=\Lambda_{t-1}^{\mathrm{inv}}$
16:     end if
17:     if $m<n$ then
18:        $W_{t+1}=W_{t}-\eta_{t}\Lambda_{t}^{\mathrm{inv}}M_{t}$
19:     else
20:        $W_{t+1}=W_{t}-\eta_{t}M_{t}\Lambda_{t}^{\mathrm{inv}}$
21:     end if
22:  end for
(i) A key modification is the adaptive selection of the preconditioning side. Instead of fixing the preconditioner to one side, we dynamically choose to precondition on the side corresponding to the smaller dimension of the gradient matrix $G_{t}$ in $\mathbb{R}^{D_{1}\times D_{2}}$. Specifically, if $D_{1}<D_{2}$, the preconditioner is formed from $G_{t}G_{t}^{T}$ and applied from the left; otherwise (if $D_{2}<D_{1}$), it is formed from $G_{t}^{T}G_{t}$ and applied from the right. This strategy, which aims to minimize computational overhead associated with the preconditioner, is also used by the Muon Optimizer. For 1D parameters (i.e., gradient vectors g), where one dimension is 1, this "smaller dimension" rule naturally leads to a scalar preconditioner (derived from $g^{T}g$). This effectively implements a form of AdaGrad-Norm [Streeter and McMahan, 2010, Ward et al., 2020] for these vector parameters. Consequently, this unified approach allows ASGO to efficiently handle all parameter structures, including 1D vectors, without needing a separate optimizer (like AdamW [Loshchilov and Hutter, 2017]) that is often employed for such parameters in methods like Muon [Jordan et al., 2024]. The convergence guarantees can be extended to this adaptive side selection, as each choice (left or right preconditioning) maintains a valid ASGO-like update structure (operating on $W$ or $W^{T}$, respectively) [^3] For instance, applying ASGO with right-side preconditioning to $W$ is equivalent to applying the original left-side ASGO formulation to $W^{T}$..
(ii) Exponential moving averages are used to update the preconditioner and incorporate momentum, following established practice, which results in Adam’s improvements across various tasks over AdaGrad.
(iii) The preconditioner matrix is inverted every $\tau$ iterations, as is done in distributed Shampoo [Shi et al., 2023] and SOAP [Vyas et al., 2024]. This modification increases memory requirements but substantially reduces computational overhead while improving algorithmic stability.
Special Design for Transformers.
Furthermore, we introduce a specialized adaptation of ASGO for query and key matrices in attention layers. Recent work by Zhang et al. [2024b] has demonstrated that the Hessian structure of query and key layers differs significantly from conventional MLP layers. As illustrated in Figure 3, the number of dense blocks corresponds to the number of attention heads rather than output neurons. This observation aligns with the forward computation of multi-head attention, where attention scores are computed independently across different subspaces, suggesting that query and key parameters could be optimized in a head-wise manner. To leverage this insight, we reshape query and key parameters from matrices in $R^{n\times hd}$ to three-dimensional tensors $R^{h\times n\times d}$, and apply our optimization algorithm independently to each head’s subspace. This restructuring reduces both memory consumption and computational complexity by decreasing the matrix size from $O(h^{2}d^{2})$ into $O(hd^{2})$. For instance, in NanoGPT [Karpathy, 2022], this adjustment reduces the preconditioning size of a single query/key parameter from approximately $10^{6}$ elements into $10^{5}$. The empirical performance of this modification is evaluated in Section C.
A Diagonal Variant of ASGO.
Drawing inspiration from Duchi et al. [2011] , we also implement a variant of ASGO using diagonal matrices, which we denote as DASGO, presented in Algorithm 3. DASGO can be viewed as a lightweight version of ASGO, which eliminates the need to compute the inverse square root of full matrices, and reduces memory requirements to a level comparable with Adam-mini [Zhang et al., 2024b]. It also makes the choice of which side of $M_{t}$ to precondition unimportant in terms of computational effort, in contrast to ASGO. For DASGO, we choose to apply the diagonal preconditioner on the right side, aligning with the neuron architecture in DNNs. However, since DASGO only employs a diagonal preconditioner, it fails to recover the superior theoretical properties of ASGO under block-wise diagonal Hessian settings . We further empirically examine this tradeoff in Section 7.
@@
1:  Input: $W_{0}\in\mathbb{R}^{m\times n}$, $\epsilon\in\mathbb{R}$, lr schedule $\{\eta_{t}\}$, momentums $\beta_{1},\beta_{2}\in[0,1)$, and batch size $M\in\mathbb{N}$
@@
5:     $G_{t}=\frac{1}{M}\sum_{\xi\in\mathcal{B}_{t}}\nabla_{W}f(W_{t};\xi)$
6:     $M_{t}=\beta_{1}M_{t-1}+(1-\beta_{1})G_{t}$
7:     $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})\operatorname{diag}(G_{t}^{\top}G_{t}$)
8:     $W_{t+1}=W_{t}-\eta_{t}M_{t}\operatorname{diag}(v_{t}+\epsilon)^{-\frac{1}{2}}$
@@
Appendix C Details of Empirical Experiments
C.1 Pretraining NanoGPT
Experimental Setup:
As an initial experiment, we compared the performance of ASGO, DASGO, Muon, Shampoo, and Adam-W for training NanoGPT, which consists of 6 Transformer layers, 6 attention heads, and an embedding dimension of 384, on the Shakespeare character-level dataset with a sequence length of 256 tokens. For all algorithms, we trained the model for 20 epochs, where each epoch contained 128 steps, using a batch size 128 and the OneCycle learning rate (lr) schedule.
Before discussing hyperparameter (HP) tuning, we note that recent research [Shi et al., 2023, Morwani et al., 2024, Lin et al., 2024] has suggested treating the inverse order (IO) in Shampoo’s preconditioner as a tunable HP rather than using the standard value of $-1/4$. Notably, Morwani et al. [2024] demonstrated that Shampoo with IO$=-1/2$ provides a superior approximation of the full-matrix AdaGrad optimizer. Consequently, we treated the IO as a Shampoo HP in our pretraining experiments. Additionally, Shampoo’s performance is highly dependent on the initialization of its preconditioner matrices, as this is crucial for accurately approximating the Empirical Fisher Information Matrix [Morwani et al., 2024]. To address this initialization sensitivity, we implemented a preconditioner warmup phase specifically for Shampoo. Following the methodology in Ren and Goldfarb [2021], we dedicated the first epoch exclusively to accumulating statistics for the preconditioner matrices without updating model parameters. This approach yields a more robust estimation of the preconditioner, which significantly improves the stability of Shampoo during subsequent training iterations. Figure 5 depicts the training loss and validation loss for training NanoGPT with Shampoo, DASGO, Muon, ASGO and Adam-W.
Hyperparameter Tuning
@@ C.1 Pretraining NanoGPT
To ensure optimal performance for each optimizer, we conducted a comprehensive HP search using random search [Bergstra and Bengio, 2012, Choi et al., 2019] within the following predefined ranges: initial lr: $[10^{-5},10^{-1}]$; $\beta_{1}$: $[0.7,0.99]$; $\beta_{2}$: $[0.7,0.99]$; and selected the warmup factor from the set $[0.1,0.15,0.2,0.25,0.3]$. For the optimizers Shampoo and ASGO, we selected the update frequency $\tau$ from the set $[5,10,15]$ and for Shampoo, we selected its IO as either $-\frac{1}{2}$ or $-\frac{1}{4}$. HPs were selected based on validation loss after 20 epochs of training. For each optimizer, we performed a random search for 16 hours to find the optimal HP combination. These best HP choices for training NanoGPT are given in Table 3.
Table 3: Hyperparameter selection for NanoGPT experiment